If a car at rest accelerates uniform to a speed of 144 km per hour in 20 seconds then it cover a distance of

Grade:11

1 Answers

Pranay
6 years ago
u=0 m/s , v=40 m/s t=20 sec. a=(v-u)/t= 2 m/s.Now s=ut+1/2at.t.Put the values and you will get it as-20 m. And.